OpenStudy (anonymous):

which value of x makes the open sentence true 10-x=3x-4

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

10-x=3x-4 +x +x = 10 = 4x-4 +4 +4 = 14 = 4x / / 4 4 = 3.5 = x :)